超伝導物質科学 2 単位 (選択)

Superconductivity and superconducting materials  岸本豊 · 教授/ システム創生工学専攻光システム工学コース光機能材料工学講座

【授業目的】超伝導現象の基礎および最近の進展について修得させる．【授業概要】量子力学の基礎を復習し，超伝導現象の基礎ならびに最近の進展に

ついて講義する．この授業は工業に関する科目である．【授業形式】講義【キーワード】超伝導，クーパー対，エネルギーギャップ，状態密度，異方的超

伝導【先行科目】 [先行科目] 【関連科目】 [関連科目] 【履修要件】 [要件] 【履修上の注意】 [注意] 【到達目標】超伝導の基礎概念を理解する．【授業計画】

1. はじめに

2. 量子力学の復習 (1) シュレディンガ-方程式と波動関数

3. 量子力学の復習 (2) 演算子と期待値

4. 量子力学の復習 (3) 摂動論

5. 金属の自由電子モデル

6. 超伝導の現象論

7. マイスナー効果の量子力学的説明

8. クーパ対の形成

9. エネルギーギャップの形成

10. 状態密度および有限温度での取り扱い

11. 測定技術-NMR の紹介-

12. 最近の超伝導体から (1) 強結合超伝導体

13. 最近の超伝導体から (2) 重い電子系超伝導体

14. 最近の超伝導体から (3) 銅酸化物高温超伝導体

15. まとめと将来の展望【成績評価基準】講義時に課すレポートで評価する．【教科書】講義時に紹介する．【参考書】講義時に紹介する．【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216770 【対象学生】 [対象学生]

{【連絡先】 \atop ⇒ 岸本豊 (A202)}

Superconductivity and superconducting materials 2 units (selection)

Yutaka Kishimoto · Professor / Optical Materials and Devices, Optical Systems Engineering, Systems Innovation Engineering

Target⟩ To understand basics and recent developments in superconductivity. Outline⟩ This class reviews quantum mechanics and introduces superconductivity

and recent superconducting materials. Style⟩ Lecture �������⟩ superconductivity, Cooper pair, superconducting energy gap, density

of states, anisotropic superconductivity Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Requirement⟩ [要件] Notice⟩ [注意] Goal⟩ To understand the outline of superconductivity.

Schedule⟩

1. Introduction to superconductivity

2. Review of quantum mechanics (1), Schrödinger equation and wave function

3. Review of quantum mechanics (2), operators and expected values

4. Review of quantum mechanics (3), perturbation theory

5. Free electron model of metals

6. Phenomenological theory of superconductivity

7. To understand Meissner effect on the basis of quantum mechanics

8. Formation of Cooper pair

9. Formation of superconducting energy gap

10. Density of states, and energy gap at finite temperatures

11. Introduction to Nuclear Magnetic Resonance method

12. Frontier of superconductivity (1), strong coupling superconductor

13. Frontier of superconductivity (2), heavy Fermion superconductor

14. Frontier of superconductivity (3), copper oxide high Tc superconductor

15. Summary Evaluation Criteria⟩ Reports on several subjects in the class. Textbook⟩ To be introduced in the class. Reference⟩ To be introduced in the class.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216770 Student⟩ [対象学生] Contact⟩ ⇒ Yutaka Kishimoto (A202)